OpenStudy (anonymous):

Use the following equation to answer the questions below. 3x^2 + 6x = 12 a=___ b=____ c=____

OpenStudy (anonymous):

a = 3 b = 6 c = -12

jimthompson5910 (jim_thompson5910):

Here's a similar example: 5x^2 + 7x = 9 5x^2 + 7x - 9 = 0 5x^2 + 7x + (-9) = 0 The equation above is in the form ax^2+bx+c=0 where a = 5, b = 7 and c = -9 So... a=__5__ b=__7__ c=__-9__ Note: this is NOT the answer as it's an answer to a similar problem.

OpenStudy (anonymous):

thanks!